# Navigation World Models

### v1

@@ 3.3 Navigation Planning with World Models
Formally, given the latent encoding $s_{0}$ and navigation target $s^{*}$, we look for a sequence of actions $(a_{0},...,a_{T})$ that maximizes the likelihood of reaching $s^{*}$. Let $\mathcal{S}({s}_{T},s^{*})$ represent the unnormalized score for reaching state $s^{*}$ with $s_{T}$ given the initial condition $s_{0}$, actions $\mathbf{a}=(a_{0},\dots,a_{T})$, and states $\mathbf{s}=({s}_{1},\dots s_{T})$ obtained by autoregressively rolling out the NWM: $\mathbf{s}\sim F_{\theta}(\cdot|s_{0},\mathbf{a})$.
We define the energy function $\mathcal{E}(s_{0},a_{0},\dots,a_{T},s_{T})$, such that minimizing the energy corresponds to maximizing the unnormalized perceptual similarity score and following potential constraints on the states and actions:
$\mathcal{E}(s_{0},a_{0},\dots,a_{T},s_{T})=-\mathcal{S}(s_{T},s^{*})+\sum_{\tau=0}^{T}\mathbb{I}(a_{\tau}\notin\mathcal{A}_{\text{valid}})+\sum_{\tau=0}^{T}\mathbb{I}(s_{\tau}\notin\mathcal{S}_{\text{safe}}),$ (4)
The similarity is computed by decoding $s^{*}$ and $s_{T}$ to pixels using a pretrained VAE decoder (Blattmann et al., 2023) and then measuring the perceptual similarity (Zhang et al., 2018a; Fu et al., 2024). Constraints like “never go left then right” can be encoded by constraining $a_{\tau}$ to be in a valid action set $\mathcal{A}_{\text{valid}}$, and “never explore the edge of the cliff” by ensuring such states $s_{\tau}$ are in $\mathcal{S}_{\text{safe}}$. $\mathbb{I}(\cdot)$ denotes the indicator function that applies a large penalty if any action or state constraint is violated.
@@
$a_{0},\dots,a_{T}=\arg\min_{a_{0},\dots,a_{T}}\mathbb{E}_{\mathbf{s}}\left[\mathcal{E}(s_{0},a_{0},\dots,a_{T},s_{T})\right]$ (5)
This objective can be reformulated as a Model Predictive Control (MPC) problem, and we optimize it using the Cross-Entropy Method (Rubinstein, 1997), a simple derivative-free and population-based optimization method which was recently used with with world models for planning (Zhou et al., 2024; Hansen et al., ). We include an overview of the Cross Entropy Method and the full optimization technical details in Appendix 7.
Ranking Navigation Trajectories. Assuming we have an existing navigation policy $\Pi(\mathbf{a}|s_{0},s^{*})$, we can use NWMs to rank sampled trajectories. Here we use NoMaD (Sridhar et al., 2024), a state-of-the-art navigation policy for robotic navigation. To rank trajectories, we draw multiple samples from $\Pi$ and choose the one with the lowest energy, like in Eq. 5.
@@ 4.1 Experimental Setting
Datasets. For all robotics datasets (SCAND (Karnan et al., 2022), TartanDrive (Triest et al., 2022), RECON (Shah et al., 2021), and HuRoN (Hirose et al., 2023)), we have access to the location and rotation of robots, allowing us to infer relative actions compare to current location (see Eq. 2). To standardize the step size across agents, we divide the distance agents travel between frames by their average step size in meters, ensuring the action space is similar for different agents, additionally we filter out backward movements, following NoMaD (Sridhar et al., 2024). Additionally, we use unlabeled Ego4D (Grauman et al., 2022) videos, where the only action we consider is time shift. SCAND provides video footage of socially compliant navigation in diverse environments, TartanDrive focuses on off-road driving, RECON covers open-world navigation, HuRoN captures social interactions. We train on unlabeled Ego4D videos and GO Stanford (Hirose et al., 2018) serves as an unknown evaluation environment. For the full details, see Appendix 8.1.
Evaluation Metrics. We evaluate predicted navigation trajectories using Absolute Trajectory Error (ATE) for accuracy and Relative Pose Error (RPE) for pose consistency (Sturm et al., 2012). To check how semantically similar are world model predictions to ground truth images, we apply LPIPS (Zhang et al., 2018b) and DreamSim (Fu et al., 2024), measuring perceptual similarity by comparing deep features, and PSNR for pixel-level quality. For image and video synthesis quality, we use FID (Heusel et al., 2017) and FVD (Unterthiner et al., 2019) which evaluate the generated data distribution. See Appendix 8.1 for more details.
@@
  • DIAMOND (Alonso et al., ) is a diffusion world model based on the UNet (Ronneberger et al., 2015) architecture. We use DIAMOND in the offline-reinforcement learning setting following their public code. The diffusion model is trained to autoregressively predict at $56$x56565656 resolution alongside an upsampler to get to $224$x224224224224 resolution predictions. We alter their action embedding to use a linear layer to work with our continuous actions.
  • GNM (Shah et al., 2023) is a general goal-conditioned navigation policy trained on a dataset soup of robotic navigation datasets with a fully connected trajectory prediction network. GNM is trained on multiple datasets including SCAND, TartanDrive, GO Stanford, and RECON.
  • NoMaD (Sridhar et al., 2024) extends GNM using a diffusion policy for predicting trajectories for robot exploration and visual navigation. NoMaD is trained on the same datasets used by GNM and on HuRoN.
Implementation Details. In the default experimental setting we use a CDiT-XL of $1B$ parameters with context of $4$ frames, a total batch size of $1024$, and $4$ different navigation goals, leading to a final total batch size of $4096$. We use the Stable Diffusion (Blattmann et al., 2023) VAE tokenizer, similar as in DiT (Peebles and Xie, 2023). We use the AdamW (Loshchilov, 2017) optimizer with a learning rate of $58e-$ 5. After training, we sample $5$ times from each model to report mean and std results. XL sized model are trained on $8$ H100 machines, each with $8$ GPUs. Unless otherwise mentioned, we use the same setting as in DiT-*/2 models.
@@ 4.2 Ablations
Model Size and CDiT. We compare CDiT (see Section 3.2), compared to a standard DiT in which all context tokens are fed as inputs. We hypothesize that for navigating known environments, the capacity of the model is the most important, and the results in Figure 6, indicate that CDiT indeed performs better with models of up to $1$B parameters, while consuming less than ×$\times 2$ FLOPs. Surprisingly, even with equal amount of parameters (e.g, CDiT-L compared to DiT-XL), CDiT is ×$\times 4$ faster and performs better.
@@
Figure 7: Ranking an external policy’s trajectories using NWM. We use NoMaD (Sridhar et al., 2024) to sample multiple trajectory predictions to navigate from the observation image to the goal image. We then simulate each of these trajectories using NWM and rank them with LPIPS (Zhang et al., 2018b). With NWM we can accurately choose trajectories that are closer to the groundtruth trajectory. Click the image to play examples in a browser.
@@
Table 2: Goal Conditioned Visual Navigation. ATE and RPE results on RECON, predicting $2$ second trajectories. NWM achieves improved results on all metrics compared to previous approaches NoMaD (Sridhar et al., 2024) and GNM (Shah et al., 2023).
@@ 4.3 Video Prediction and Synthesis
Generation Quality. To evaluate video quality, we auto-regressively predict videos at $4$ FPS for $16$ seconds to create videos, while conditioning on ground truth actions. We then evaluate the quality of videos generated using FVD, compared to DIAMOND (Alonso et al., ). The results in Figure 6 indicate that NWM outputs higher quality videos.
@@ 4.4 Planning Using a Navigation World Model
Standalone Planning. We demonstrate that our World Model can be effectively used independently for goal-conditioned navigation. We condition it on past observations and a goal image, and use the Cross-Entropy Method to find a trajectory that minimizes the LPIPS similarity of the last predicted image to the goal image (see Equation 5). To rank an action sequence, we execute the NWM and measure LPIPS between the last state and the goal $3$ times to get an average score. We generate trajectories of length $8$, with temporal shift of $k=0.25$. We evaluate the model performance in Table 3. We find that using a NWM for planning leads to competitive results with state-of-the-art policies.
Figure 9: Planning with Constraints Using NWM. We visualize trajectories planned with NWM under the constraint of moving left or right initially, followed by forward motion. The planning objective is to reach the same final position and orientation as the ground truth (GT) trajectory. Shown are the costs for proposed trajectories 0, 1, and 2, with trajectory 0 (in green) achieving the lowest cost.
Planning with Constraints. Planning within a world model allows incorporating constraints. For example, an agent might be constrained to move in a straight line or take only one turn. We demonstrate that we can plan using a NWM while satisfying constraints. In forward-first, the agent moves forward for 5 steps before turning left or right for 3 steps. In left-right first, the agent turns left or right for 3 steps before moving forward for 5 steps. In straight then forward, the agent moves straight in any direction for 3 steps before moving forward. To set these constraints, we can simply hardwire the corresponding actions to be zero, for example, in left-right first, we set forward motion to zero for the first three actions and run Standalone Planning to optimize only the remaining actions. See Appendix 8.2 for the full details. Finally, we report the norm of the difference between the final position and final yaw for each setting, relative to unconstrained planning where the agent go in a straight line in any direction.
The results in Table 3 indicate that NWM can plan effectively while satisfying the constraints, observing only minor differences in the planning performance. We include examples of planned trajectories under left-right first constraint in Figure 9.
Using a Navigation World Model for Ranking. NWM can enhance existing navigation policies in goal-conditioned navigation. Conditioning NoMaD on past observations and a goal image, we sample $n\in\{16,32\}$ trajectories, each of length $8$, and evaluate them by autoregressively following the actions using NWM. Finally, we rank each trajectory’s final prediction by measuring LPIPS similarity with the goal image (see Figure 7). We report ATE and RPE on all in-domain datasets (see Table 3) and find that ranking trajectories yields state-of-the-art navigation performance, with more sampled trajectories yielding better results.
@@ 4.5 Generalization to Unknown Environments
Here we experiment with adding unlabeled data, and ask whether NWM can make predictions in new environments using imagination. Clearly, In this experiment we train a model on all in-domain datasets, as well as a susbet of unlabeled videos from Ego4D, where we only have access to the time-shift action. We train a CDiT-XL model and test it on the Go Stanford dataset as well as other random images. We report the results in Table 4, finding that training on unlabeled data leads to significantly better video predictions according to all metrics, including improved generation quality. We include qualitative examples in Figure 8. Compared to in-domain (Figure 3), the model breaks faster and expectedly hallucinates paths as it generates traversals of imagined environments.
@@ 5 Limitations
We identify multiple limitations. First, when applied to out of distribution data, we noticed that the model tends to slowly lose context and generates next states that resemble the training data, a phenomena that was observed in image generation and is known as mode collapse (Thanh-Tung and Tran, 2020; Srivastava et al., 2017). We include such an example in Figure 10. Second, while the model can plan, it struggles with simulating temporal dynamics like pedestrian motion (although in some cases it does). Both limitations are likely to be solved with longer context and more training data. Additionally, the model currently utilizes $3$ DOF navigation actions, but extending to $6$ DOF navigation and potentially more (like controlling the joints of a robotic arm) are possible as well, which we leave for future work.
6 Conclusion
Our proposed Navigation World Model (NWM) offers a scalable, data-driven approach to learning navigation policies. NWM is trained across diverse environments with our CDiT architecture, and is able to adapt flexibly to various scenarios. NWM can independently plan or rank an external policy by simulating navigation outcomes, which also allows it to incorporate new constraints. This approach bridges learning from video, visual navigation, and model-based planning and could potentially open the door to self-supervised systems that not only perceive but also act.

### v2

@@ 3.3 Navigation Planning with World Models
Formally, given the latent encoding $s_{0}$ and navigation target $s^{*}$, we look for a sequence of actions $(a_{0},...,a_{T-1})$ that maximizes the likelihood of reaching $s^{*}$. Let $\mathcal{S}({s}_{T},s^{*})$ represent the unnormalized score for reaching state $s^{*}$ with $s_{T}$ given the initial condition $s_{0}$, actions $\mathbf{a}=(a_{0},\dots,a_{T-1})$, and states $\mathbf{s}=({s}_{1},\dots s_{T})$ obtained by autoregressively rolling out the NWM: $\mathbf{s}\sim F_{\theta}(\cdot|s_{0},\mathbf{a})$.
We define the energy function $\mathcal{E}(s_{0},a_{0},\dots,a_{T-1},s_{T})$, such that minimizing the energy corresponds to maximizing the unnormalized perceptual similarity score and following potential constraints on the states and actions:
$\mathcal{E}(s_{0},a_{0},\dots,a_{T-1},s_{T})=-\mathcal{S}(s_{T},s^{*})+$ (4)
$+\sum_{\tau=0}^{T-1}\mathbb{I}(a_{\tau}\notin\mathcal{A}_{\text{valid}})+\sum_{\tau=0}^{T-1}\mathbb{I}(s_{\tau}\notin\mathcal{S}_{\text{safe}}),$
The similarity is computed by decoding $s^{*}$ and $s_{T}$ to pixels using a pretrained VAE decoder [4] and then measuring the perceptual similarity [75, 14]. Constraints like “never go left then right” can be encoded by constraining $a_{\tau}$ to be in a valid action set $\mathcal{A}_{\text{valid}}$, and “never explore the edge of the cliff” by ensuring such states $s_{\tau}$ are in $\mathcal{S}_{\text{safe}}$. $\mathbb{I}(\cdot)$ denotes the indicator function that applies a large penalty if any action or state constraint is violated.
@@
$\arg\min_{a_{0},\dots,a_{T-1}}\mathbb{E}_{\mathbf{s}}\left[\mathcal{E}(s_{0},a_{0},\dots,a_{T-1},s_{T})\right]$ (5)
This objective can be reformulated as a Model Predictive Control (MPC) problem, and we optimize it using the Cross-Entropy Method [48], a simple derivative-free and population-based optimization method which was recently used with with world models for planning [77]. We include an overview of the Cross-Entropy Method and the full optimization technical details in Appendix 7.
Ranking Navigation Trajectories. Assuming we have an existing navigation policy $\Pi(\mathbf{a}|s_{0},s^{*})$, we can use NWMs to rank sampled trajectories. Here we use NoMaD [55], a state-of-the-art navigation policy for robotic navigation. To rank trajectories, we draw multiple samples from $\Pi$ and choose the one with the lowest energy, like in Eq. 5.
@@ 4.1 Experimental Setting
Datasets. For all robotics datasets (SCAND [30], TartanDrive [60], RECON [52], and HuRoN [27]), we have access to the location and rotation of robots, allowing us to infer relative actions compare to current location (see Eq. 2). To standardize the step size across agents, we divide the distance agents travel between frames by their average step size in meters, ensuring the action space is similar for different agents. We further filter out backward movements, following NoMaD [55]. Additionally, we use unlabeled Ego4D [18] videos, where the only action we consider is time shift. SCAND provides video footage of socially compliant navigation in diverse environments, TartanDrive focuses on off-road driving, RECON covers open-world navigation, HuRoN captures social interactions. We train on unlabeled Ego4D videos and GO Stanford [24] serves as an unknown evaluation environment. For the full details, see Appendix 8.1.
Evaluation Metrics. We evaluate predicted navigation trajectories using Absolute Trajectory Error (ATE) for accuracy and Relative Pose Error (RPE) for pose consistency [57]. To check how semantically similar are world model predictions to ground truth images, we apply LPIPS [76] and DreamSim [14], measuring perceptual similarity by comparing deep features, and PSNR for pixel-level quality. For image and video synthesis quality, we use FID [23] and FVD [64] which evaluate the generated data distribution. See Appendix 8.1 for more details.
@@
  • DIAMOND [1] is a diffusion world model based on the UNet [47] architecture. We use DIAMOND in the offline-reinforcement learning setting following their public code. The diffusion model is trained to autoregressively predict at $56$x56565656 resolution alongside an upsampler to obtrain $224$x224224224224 resolution predictions. To condition on continuous actions, we use a linear embedding layer.
  • GNM [53] is a general goal-conditioned navigation policy trained on a dataset soup of robotic navigation datasets with a fully connected trajectory prediction network. GNM is trained on multiple datasets including SCAND, TartanDrive, GO Stanford, and RECON.
  • NoMaD [55] extends GNM using a diffusion policy for predicting trajectories for robot exploration and visual navigation. NoMaD is trained on the same datasets used by GNM and on HuRoN.
Implementation Details. In the default experimental setting we use a CDiT-XL of $1B$ parameters with context of $4$ frames, a total batch size of $1024$, and $4$ different navigation goals, leading to a final total batch size of $4096$. We use the Stable Diffusion [4] VAE tokenizer, similar as in DiT [44]. We use the AdamW [39] optimizer with a learning rate of $58e-$ 5. After training, we sample $5$ times from each model to report mean and std results. XL sized model are trained on $8$ H100 machines, each with $8$ GPUs. Unless otherwise mentioned, we use the same setting as in DiT-*/2 models.
@@ 4.2 Ablations
Model Size and CDiT. We compare CDiT (see Section 3.2) with a standard DiT in which all context tokens are fed as inputs. We hypothesize that for navigating known environments, the capacity of the model is the most important, and the results in Figure 6, indicate that CDiT indeed performs better with models of up to $1$B parameters, while consuming less than 2×2$\times$2 × FLOPs. Surprisingly, even with equal amount of parameters (e.g, CDiT-L compared to DiT-XL), CDiT is 4×4$\times$4 × faster and performs better.
@@
Figure 7: Ranking an external policy’s trajectories using NWM. To navigate from the observation image to the goal, we sample trajectories from NoMaD [55], simulate each of these trajectories using NWM, score them (see Equation 4), and rank them. With NWM we can accurately choose trajectories that are closer to the groundtruth trajectory. Click the image to play examples in a browser.
@@
Table 2: Goal Conditioned Visual Navigation. ATE and RPE results on RECON, predicting $2$ second trajectories. NWM achieves improved results on all metrics compared to previous approaches NoMaD [55] and GNM [53].
@@ 4.3 Video Prediction and Synthesis
Generation Quality. To evaluate video quality, we auto-regressively predict videos at $4$ FPS for $16$ seconds to create videos, while conditioning on ground truth actions. We then evaluate the quality of videos generated using FVD, compared to DIAMOND [1]. The results in Figure 6 indicate that NWM outputs higher quality videos.
@@ 4.4 Planning Using a Navigation World Model
Standalone Planning. We demonstrate that NWM can be effectively used independently for goal-conditioned navigation. We condition it on past observations and a goal image, and use the Cross-Entropy Method to find a trajectory that minimizes the LPIPS similarity of the last predicted image to the goal image (see Equation 5). To rank an action sequence, we execute the NWM and measure LPIPS between the last state and the goal $3$ times to get an average score. We generate trajectories of length $8$, with temporal shift of $k=0.25$. We evaluate the model performance in Table 3. We find that using a NWM for planning leads to competitive results with state-of-the-art policies.
Figure 9: Planning with Constraints Using NWM. We visualize trajectories planned with NWM under the constraint of moving left or right first, followed by forward motion. The planning objective is to reach the same final position and orientation as the ground truth (GT) trajectory. Shown are the costs for proposed trajectories 00, $1$, and $2$, with trajectory 00 (in green) achieving the lowest cost.
Planning with Constraints. World models allow planning under constraints—for example, requiring straight motion or a single turn. We show that NWM supports constraint-aware planning. In forward-first, the agent moves forward for 5 steps, then turns for 3. In left-right first, it turns for 3 steps before moving forward. In straight then forward, it moves straight for 3 steps, then forward. Constraints are enforced by zeroing out specific actions; e.g., in left-right first, forward motion is zeroed for the first 3 steps, and Standalone Planning optimizes the rest. We report the norm of the difference in final position and yaw relative to unconstrained planning. Results (Table 3) show NWM plans effectively under constraints, with only minor performance drops (see examples in Figure 9).
Using a Navigation World Model for Ranking. NWM can enhance existing navigation policies in a goal-conditioned navigation. Conditioning NoMaD on past observations and a goal image, we sample $n\in\{16,32\}$ trajectories, each of length $8$, and evaluate them by autoregressively following the actions using NWM. Finally, we rank each trajectory’s final prediction by measuring LPIPS similarity with the goal image (see Figure 7). We report ATE and RPE on all in-domain datasets (Table 3) and find that NWM-based trajectory ranking improves navigation performance, with more samples yielding better results.
@@ 4.5 Generalization to Unknown Environments
Here we experiment with adding unlabeled data, and ask whether NWM can make predictions in new environments using imagination. In this experiment, we train a model on all in-domain datasets, as well as a susbet of unlabeled videos from Ego4D, where we only have access to the time-shift action. We train a CDiT-XL model and test it on the Go Stanford dataset as well as other random images. We report the results in Table 4, finding that training on unlabeled data leads to significantly better video predictions according to all metrics, including improved generation quality. We include qualitative examples in Figure 8. Compared to in-domain (Figure 3), the model breaks faster and expectedly hallucinates paths as it generates traversals of imagined environments.
@@ 5 Limitations
We identify multiple limitations. First, when applied to out of distribution data, the model tends to slowly lose context and generates next states that resemble the training data, a phenomena that was observed in image generation and is known as mode collapse [58, 56]. We include such an example in Figure 10. Second, while the model can plan, it struggles with simulating temporal dynamics like pedestrian motion (although in some cases it does). Both limitations are likely to be solved with longer context and more training data. Additionally, the model currently utilizes $3$ DoF navigation actions, but extending to $6$ DoF navigation and potentially more (like controlling the joints of a robotic arm) are possible as well, which we leave for future work.
6 Discussion
Our proposed Navigation World Model (NWM) offers a scalable, data-driven approach to learning world models for visual navigation; However, we are not exactly sure yet what representations enable this, as our NWM does not explicitly utilize a structured map of the environment. One idea, is that next frame prediction from an egocentric point of view can drive the emergence of allocentric representations [65]. Ultimately, our approach bridges learning from video, visual navigation, and model-based planning and could potentially open the door to self-supervised systems that not only perceive but can also plan to inform action.